OpenStudy (anonymous):

The sum of two fractions is 11/12, their product is 1/6, what is the lesser of the two fractions?

OpenStudy (anonymous):

\[ \begin{align} a+b&={11\over12}\\ ab&={1\over6}\\ a\left({11\over12}-a\right)&={1\over6}\\ 11a-12a^2&=2\\ 12a^2-11a+2&=0\\ 12a^2-8a-3a+2&=0\\ (4a-1)(3a-2)&=0\\ \end{align}\\ \boxed{a={1\over4}}\qquad\boxed{b={3\over2}} \]

OpenStudy (aakashsudhakar):

x + y = 11/12 xy = 1/6 = 24/144 x = 8/12 = 2/3 y = 3/12 = 1/4 (3/12) or (1/4) is the lesser of the two fractions.

OpenStudy (anonymous):

Good work! Thanks to both of you